A 6.5g mass of ice with a temperature of 0 degree
Answer #1

Given

mass of the ice m = 6.5 g

temperature of lake water = 150C = 288 K;

temperature of ice = 00C = 273 K

specific heat capacity of water C = 4.18 J/g-K

the latent heat of fusion of ice L = 334 J/g

(1)

The final temperature will be 150 C as the lake has too much mass as compared to the ice and thus it's temperature change will be extremely small to be noticed.

(2)

heat absorbed by ice to melt (Q1) = m*L = 6.5*334 = 2171 J

heat absorbed by the melted ice (at 00C) to reach 150C (Q2) = m*C*ΔT = 6.5*4.18*15 =  407.55 K

Hence the net heat absorbed by the ice ΔQ = Q1 + Q2 = 2578.55 J

(3)

The water originally in the lake loses water.

(4)

Zeroth law of thermodynamics which states that when two bodies are in thermal contact with each other then the heat lost by the body at higher temperature is equal to the heat gained by the body at lower temperature.

Hence the heat lost by the lake = ΔQ = 2578.55 J (the same heat that was calculated in part 3)